Question 4

On 1 January 2015, Vivek plc issued 50,000, £100 2 per cent debentures to investors for £55 each. The debentures are redeemable at their par value of £100 in five years’ time, 31 December 2019. The interest rate implicit for the debenture is 15.62% per annum.

a) In accordance with IAS 39 Financial Instruments: Recognition and Measurement: Calculate the finance cost for Vivek plc in respect of the financial instrument, for each of the five years ended December 2019, and the liability of the financial instrument at the end of each of the year. (10 marks)
Answer #1

Answer:

As per IAS 39, Debentures are to be measured at amortized cost. Effective interest method applies to liabilities measured at amortized cost.

Issued value = 50000 / 100 * 55 = £27,500

Hence:

For Year Ended December 31, 2015:

Finance cost = 27500 * 15.62% = £4295.50

Cash paid = 50000 * 2% =£1,000

Liability at the end of each of the year = 27500 + (4295.50 - 1000) = 30,795.50

For year ended December 31, 2016:

Finance cost = 30795.50 * 15.62% = £4811

Cash paid = 50000 * 2% = £1,000

Liability at the end of each of the year = 30795.50 + (4811 - 1000) = 34607

Similarly we calculate for each year for next 3 years.

The Finance cost and Liability at the end of each year for Years from 2015 to 2019 are given below:

Liability balance at the Year end Amount Cash paid Finance Cost End of Year Amoritized 1,000 £ 4,296 £ 3,296 £ 30,796 1. 3,81